### 01 Apr 2026 v1

@@ 1 Introduction
A practical challenge, however, is ensuring that all students have the opportunity to engage with AI education during classroom instruction. Formal opportunities to learn about AI are limited for students, and when they occur, are often delivered as stand-alone workshops or other add-on experiences that reach a subset of students, rather than being interwoven with content area coursework that draws broad interest from the general population of students [9, 10]. Related work suggests that one promising direction is to integrate AI- and CT-related learning into middle school science in ways that connect to disciplinary content and regular classroom instruction, rather than positioning these topics only as stand-alone experiences [2, 5]. Further, designing for broad participation means developing learning experiences that are feasible within classroom constraints, reflect teacher and community input, and account for differences in local resources, available instructional time, and teacher support and preparation that can affect curricular accessibility across school contexts [6]. Logistical considerations also intersect with student-level barriers to engaging with AI education. Introductory AI materials reach students with widely different prior experiences, yet many still presume some familiarity with programming or advanced mathematics [10]. As a result, participation may be narrowed toward students who already feel positively about AI or are more likely to opt in.
@@ 3.2 Application and supporting activities
Following the four initial BFS activities in I-SAIL on Days 2 and 3, Day 4 shifted to an unplugged, paper-based tracing task that gave students additional practice carrying out the procedure by hand without the environment’s supports. Day 5 focused on unplugged science activities connecting virus transmission to AI problem-solving over contact networks. In a hands-on "Glo Germ" simulation, a small amount of fluorescent powder was placed on a student’s hands to represent an infectious agent. As students interacted, the powder spread through contact and was later revealed under UV light, helping students see how transmission moves through chains of contact. Building on this activity, a card-based task modeled the class as a contact network. Each student received a card, and connections were drawn to represent recent contacts. Green and red dot markers indicated infection states, and students discussed infection likelihood in terms of network distance, or steps, between individuals. This connected the science activity back to BFS as a procedure for tracing shortest contact paths. On Day 6, students used a simulation in I-SAIL to connect BFS reasoning to scientific problem-solving by generating and testing hypotheses about how factors such as outdoor temperature and network density affect patterns of virus spread in connected populations. On the final day, the module concluded with a career session discussing different career pathways, how AI and computing can be used in those roles, and how those uses can affect people and communities.
@@ 4 Methods
The study took place in Fall 2025 in a rural public middle school in the United States and included six classrooms across three grade levels: two 6th-grade classrooms (typically ages 11–12), two 7th-grade classrooms (typically ages 12–13), and two 8th-grade classrooms (typically ages 13–14). The module was delivered across seven days in a health science course that met in 45-minute class periods. Approximately, 30% of students at the participating middle school in each grade level were enrolled in the course, with all sections taught by the same instructor. This study was approved by North Carolina State University’s Institutional Review Board, and parental consent was obtained prior to data analysis. In total, 103 students were enrolled across the six classrooms and participated in the module as part of regular instruction. Data were collected from three primary sources: (1) pre/post knowledge assessments, (2) student work artifacts, and (3) a teacher interview. Analysis reported in this paper include only students with consent and matched data for the relevant measures; after filtering, $N=59$ students with parental consent remained ($n_{6}=14$, $n_{7}=29$, $n_{8}=16$).
Students completed matched BFS knowledge assessments immediately before Activity 1 and after Activity 4 in I-SAIL. The assessment included five multiple-choice items (K1–K5) targeting both general search intuitions and BFS procedural understanding. Each item was scored for correctness, and the assessment score was computed as the proportion correct across the five items.
Following the learning activities on Day 2 and 3, students completed a paper-based BFS tracing worksheet in class. Students were given a graph with designated start and goal nodes and were asked to carry out BFS by hand, recording (1) an “explored” list (the nodes they had already checked) and (2) a “path” list (the candidate routes they were considering as the search progressed). This was students’ first time practicing BFS without the environment’s visual supports and served both as reinforcement and as an artifact for examining their procedural understanding of BFS. After the classroom implementation, we conducted a semi-structured interview with the instructor to gather feedback on the curriculum, their experience leading the activities, and opportunities for improvement.
@@ 6 Discussion
The findings also suggest that grounding AI problem-solving in familiar contexts supported students’ understanding of BFS. By introducing the procedure through GPS-style navigation and then revisiting it through virus transmission and contact tracing, the curriculum helped connect an abstract search process to settings students could readily interpret. Teacher feedback further suggests that this subject-area connection helped students see the relevance and purpose of the procedure within the health science unit. Future work should examine how this kind of contextual grounding can be adapted across disciplines and whether encountering AI problem-solving in multiple familiar contexts supports broader understanding and use of these ideas across domains.

### 15 Apr 2026 v2

@@ Abstract
Preprint. Accepted to the 27th International Conference on Artificial Intelligence in Education (AIED 2026).
@@ 1 Introduction
A practical challenge, however, is ensuring that all students have the opportunity to engage with AI education during classroom instruction. Formal opportunities to learn about AI are limited for students, and when they occur, are often delivered as stand-alone workshops or other add-on experiences that reach a subset of students, rather than being interwoven with content area coursework that draws broad interest from the general population of students [9, 10]. Related work suggests that one promising direction is to integrate AI- and computational thinking-related topics into K–12 science instruction in ways that connect to disciplinary content and classroom curriculum, rather than positioning these topics only as stand-alone experiences [2, 5]. Further, designing for broad participation means developing learning experiences that are feasible within classroom constraints, reflect teacher and community input, and account for differences in local resources, available instructional time, and teacher support and preparation that can affect curricular accessibility across school contexts [6]. Logistical considerations also intersect with student-level barriers to engaging with AI education. Introductory AI materials reach students with widely different prior experiences, yet many still presume some familiarity with programming or advanced mathematics [10]. As a result, participation may be narrowed toward students who already feel positively about AI or are more likely to opt in.
@@ 3.2 Application and supporting activities
Following the four initial BFS activities in I-SAIL on Days 2 and 3, Day 4 shifted to an unplugged, paper-based tracing task that gave students additional practice carrying out the procedure by hand. Day 5 focused on unplugged science activities connecting virus transmission to AI problem-solving over contact networks. In a hands-on "Glo Germ" simulation, a small amount of fluorescent powder was placed on a student’s hands to represent an infectious agent. As students interacted, the powder spread through contact and was later revealed under UV light, helping students see how transmission moves through chains of contact. Building on this activity, a card-based task modeled the class as a contact network. Each student received a card, and connections were drawn to represent recent contacts. Green and red dot markers indicated infection states, and students discussed infection likelihood in terms of network distance, or steps, between individuals. This connected the science activity back to BFS as a procedure for tracing shortest contact paths. On Day 6, students used a simulation in I-SAIL to connect BFS reasoning to scientific problem-solving by generating and testing hypotheses about how factors such as temperature and network density affect patterns of virus spread in connected populations. On the final day, the module concluded with a session discussing career pathways, how AI and computing are used in those roles, and their effects on people and communities.
@@ 4 Methods
The study took place in Fall 2025 in a rural public middle school in the United States (typically ages 11–14). The module was implemented in six health science classrooms across three grade levels: two 6th-grade classrooms, two 7th-grade classrooms, and two 8th-grade classrooms. Each class met in 45-minute class periods, and the module was delivered over seven days by the same instructor. Approximately, 30% of students at the participating middle school in each grade level were enrolled in the course. In total, 103 students participated as part of regular instruction. This study was approved by North Carolina State University’s Institutional Review Board, and parental consent was obtained prior to analysis. The analyses reported include only students with consent and matched data for the relevant measures; after filtering, $N=59$ students with parental consent remained ($n_{6}=14$, $n_{7}=29$, $n_{8}=16$).
Data were collected from three primary sources: (1) pre/post knowledge assessments, (2) student work artifacts, and (3) a teacher interview. Students completed matched BFS knowledge assessments immediately before Activity 1 and after Activity 4 in I-SAIL. The assessment included five multiple-choice items (K1–K5) targeting both general search intuitions and BFS procedural understanding. Each item was scored for correctness, and the assessment score was computed as the proportion correct across the five items.
Following the learning activities on Day 2 and 3, students completed a paper-based BFS tracing worksheet. Students were given a graph with designated start and goal nodes and were asked to carry out BFS by hand, recording (1) an “explored” list (the nodes they had already checked) and (2) a “path” list (the candidate routes they were considering as the search progressed). This was students’ first time practicing BFS without the environment’s visual supports and served both as reinforcement and as an artifact for examining their procedural understanding of BFS. After the classroom implementation, we conducted a semi-structured interview with the instructor to gather feedback on the curriculum, their experience leading the activities, and opportunities for improvement.
@@ 6 Discussion
The findings also suggest that grounding AI problem-solving in familiar contexts supported students’ understanding of BFS. By introducing the procedure through GPS-style navigation and revisiting it through virus transmission and contact tracing, the curriculum connected an abstract search process to settings students could readily interpret. Teacher feedback further suggests that this subject-area connection helped students see the relevance and purpose of the procedure within the health science unit. Future work could examine how similar subject-area grounded approaches help students connect AI problem-solving to topics across their science curriculum.
@@ 7 Conclusion
{credits}
7.0.1 Acknowledgements
This research was supported by the U.S. National Science Foundation (NSF) under Grant #2405862. Any opinions, findings, and conclusions expressed in this material are those of the authors and do not necessarily reflect views of the NSF.